OpenStudy (anonymous):

198q3r2 – 184q2r2 + 18qr2

OpenStudy (cwrw238):

first find the GCF (greatest common factor) GCF of 198, 18 and 18 is §8 can u find GCF of q,q3 and q2 and of r2, r2 and r2 - thats easy

OpenStudy (cwrw238):

* second line - i meant 18

OpenStudy (anonymous):

6? I'm trying to help my bf with his math work.

OpenStudy (cwrw238):

sorry - i had the numbers wrong u need GCF of 198, 184 and 18

OpenStudy (cwrw238):

best to factor numbers like these to find GCF 198 = 2 * 99 184 = 2*2 *2 * 23 18 = 2*3*3 common factor is 2 so 2 is the GCF

OpenStudy (anonymous):

K. So 2 ist he GCF, now what?

OpenStudy (cwrw238):

now find gcf of q and r terms

OpenStudy (anonymous):

Okay. Q=1 R=2?

OpenStudy (cwrw238):

q terms - you have q, q^2 and q^3 - which of these will divide exactly into each one?

OpenStudy (anonymous):

1?

OpenStudy (cwrw238):

the gcfs are q and r^2

OpenStudy (cwrw238):

not 1 - the gcf is q - it will dviide exactly into q q2 and q3

OpenStudy (anonymous):

My bad. I was saying 1 because there was only a q there..

OpenStudy (cwrw238):

ok - so the 'complete' gcf is 2qr^2

OpenStudy (anonymous):

Yes. That's what I was thinking man, thanks alot.

OpenStudy (cwrw238):

now you divide the gcf into each of the original terms to get what goes in the brackets i[ll start you off: = 2qr^2(99q^2 - .. (because 198q^3r^2 / 2qr^2 = 99q^2)